A small block of wood of density 0.4x10^3 kg/m^3 is submerged in water at a depth of 2.9m. find
 
 a) acceleration of the block towards the surface when the block is released
b) the time for the block to reach the surdace. ignore viscosity

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
A ball whose density 0.4 × 10³ kg/m³ falls into water . Before falling into the water, velocity of ball at the surface of water , u = √{2gh}

Now, use formula for finding acceleration of ball,
a = apparent weight/mass of ball
Apparent weight = V(ρ - σ)g
Where , V is the volume of ball , ρ is density of ball and σ is density of water.
Now, apprent weight = V(0.4 - 1) × 10³ × g = -0.6gV × 10³
∴ a = -0.6gV × 10³/ρV
= -0.6g/0.4 = -3g/2 m/s²